OpenStudy (nuccioreggie):

Which number is the largest? 1.2 × 104, 4.5 × 10−3, 3.8 × 103, 7.5 × 103 1.2 × 104 4.5 × 10−3 3.8 × 103 7.5 × 103

OpenStudy (nuccioreggie):

@LegendarySadist. @mathstudent55

OpenStudy (nuccioreggie):

I got C for the highest power

OpenStudy (nuccioreggie):

@mathstudent55. Can you help me please

OpenStudy (anonymous):

\[\large 10^{4} \geq 10^{3}\]

OpenStudy (nuccioreggie):

A?

OpenStudy (anonymous):

Right. \[\large 1 \times 10^{4} \geq 9.999 \times 10^{3}\]

OpenStudy (nuccioreggie):

Ok thank u A is the answer